OpenStudy (anonymous):

giving medal If x = 40, solve for y. y = 3/8x A.15 B.8 C.64/3 D.3/8

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

substitude 40 for x in your equation. what you will get?

OpenStudy (anonymous):

i got 15 i don't think that's right

OpenStudy (anonymous):

or is it?

OpenStudy (anonymous):

3/8*40=3*5=15

OpenStudy (anonymous):

oh, ok thx

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thnx